Problem #10 (TVM - N): Toms accountant guesstimates sales growth to be 5%. How long would it take sales to double?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value($2x)
P=present value($x say)
r=rate of interest
n=time period.

2x=x*(1.05)^n

2=(1.05)^n

Taking log on both sides;

log 2=n*log (1.05)

n=log 2/log (1.05)

=14.21 years(Approx).